OpenStudy (anonymous):

For\[\int_{-\infty}^{\infty} e^{tx} \frac{1}{\sigma \sqrt{2\pi}} e^{\frac{-(x-\mu)^2}{2\sigma^2}}dx\]will we get \[e^{t \mu + \frac{1}{2}\sigma^2 t^2}\]?

OpenStudy (dan815):

the small part is hard to see

OpenStudy (anonymous):

Hmm power of e?

OpenStudy (dan815):

ya

OpenStudy (dan815):

have u learnt integration by parts yet

OpenStudy (anonymous):

\[-\frac{(x-\mu)^2}{2\sigma^2}\]

OpenStudy (anonymous):

Yes.

OpenStudy (dan815):

well before you try by parts just expand that bracket -(x-u)^2 and see if u can simply it down

OpenStudy (anonymous):

Hmm.. How are you going to "simplify it"?

OpenStudy (dan815):

like u know that e^(x+a) = same as e^x * e^a

OpenStudy (anonymous):

\[\int_{-\infty}^{\infty} e^{tx} \frac{1}{\sigma \sqrt{2\pi}} e^{\frac{-(x-\mu)^2}{2\sigma^2}}dx\]\[=\int_{-\infty}^{\infty} \frac{1}{\sigma \sqrt{2\pi}} e^{\frac{-(x-\mu)^2}{2\sigma^2}+tx}dx\]\[=\int_{-\infty}^{\infty} \frac{1}{\sigma \sqrt{2\pi}} e^{\frac{-(x-\mu)^2+2\sigma^2tx}{2\sigma^2}}dx\]\[\large =\int_{-\infty}^{\infty} \frac{1}{\sigma \sqrt{2\pi}} e^{\frac{-(x^2-2\mu x +\mu^2-2\sigma^2tx)}{2\sigma^2}}dx\]

OpenStudy (dan815):

so